# Hamiltonian and Liouvillian learning in weakly-dissipative quantum many-body systems

### v1

@@ III.1 Hamiltonian learning in the presence of weak dissipation
We assume that the analog quantum simulator to be studied is governed by a master equation with the following model Liouvillian
$\mathcal{L}(\varrho)=-i[H,\varrho]+\sum_{\eta\in\{+,-,z\}}\frac{\gamma_{\eta}}{2}\sum_{k=1}^{N}\quantity([\sigma_{k}^{\eta}\varrho,\sigma_{k}^{\eta,\dagger}]+\mathrm{h.c.}).$ (21)
@@
Figure 2: Hamiltonian learning in presence of weak dissipation. (a) The learning error $\lambda_{1}/\lambda_{2}$ as a function of the number of runs, $N_{\rm runs}$, of the simulator for different ansätze described in the main text. The colored triangles indicate the asymptotic values for $N_{\rm runs}\to\infty$. (b) Snapshots of the (normalized) reconstructed Hamiltonian parameters for some of the ansätze described in the main text. The point at which these snapshots are taken in terms of $N_{\mathrm{runs}}$ is also indicated by asterisks in the left panel. All error bars are computed via bootstrapping as explained in Appendix C, using $80$ samples. Model parameters: The Hamiltonian parameters in Eq. (21) are chosen as $B_{x}=4/5B_{z}$. Moreover, the coupling strengths are described by $4$-th order polynomials of the form $J_{i,i+1}^{z}=B_{z}\sum_{l=1}^{5}a_{l}\cdot x_{i,i+1}^{l-1}$, and $J_{i,i+2}^{z}=B_{z}\sum_{l=1}^{5}b_{l}\cdot x_{i,i+2}^{l-1}$ where $x_{i,j}=[(i+j)-(N+1)]/N$. We choose the following coefficients: $\bm{a}=(6/5,1/20,1/5,0,-2/5)$ and $\bm{b}=(1/5,1/20,-2/5,0,4/5)$. The dissipation rates are chosen such that $(\gamma_{+},\gamma_{-},\gamma_{z})=(1,1.5,2)\times 10^{-2}B^{z}\ll\|H\|$. To estimate the integrals in Eq. (8), we use $B_{z}\Delta_{t}=1/64$, i.e., $64$ equidistant times, where at each time we spend only $1/64$ of the measurements per basis, compared to the quench times $B_{z}T\in[0,0.5,1]$.
@@
III.1.1 Identifying the dominant terms of the Hamiltonian
@@
leads to a much smaller value of $\lambda_{1}/\lambda_{2}$. We emphasize that learning with the ansatz $A_{2}$ only requires measurements of $\sigma_{x}^{\otimes N}$ and $\sigma_{z}^{\otimes N}$. Therefore, some of the measurements performed for $A_{1}$ can be recycled for learning the parameters of $A_{2}$. As shown in Fig. 2a (orange line), $A_{2}$ leads to a plateau of $\lambda_{1}/\lambda_{2}$ beginning at $N_{\mathrm{runs}}\approx 3\times 10^{5}$. At this point we have identified all the dominant terms in $H$. To continue the learning process we need to extend the operator ansatz. Therefore, in the next step we want to learn subdominant terms of $H$, as well as the Liouvillian.
III.1.2 Learning the sub-dominant terms of the Hamiltonian, and learning the Liouvillian
@@
III.1.3 Learning the spatial variations of the couplings
As can be seen in Fig. 2a (red line), the plateau in $A_{4}$ is reached at $N_{\mathrm{runs}}\approx 10^{7}$, which again exceeds our available measurement budget. Therefore, we are not allowed to conclude that ansatz $A_{4}$ is insufficient to describe the Hamiltonian in Eq. (II.1) by only considering the ratio $\lambda_{1}/\lambda_{2}$. However, at this point, one can choose to test other ansätze and compare their corresponding reconstructed parameters, or to see if smaller values of $\lambda_{1}/\lambda_{2}$ are achieved. As one typically expects small spatial variations in the couplings of $H$, and $A_{4}$ was already a good approximation of $H$, we want to test for spatial variations on top of the spatially homogeneous ansatz $A_{4}$. To this end, we simply treat the parametrization of the ansatz $A_{4}$ as a regularization of the cost function as explained in Section II.6. That is, we start with the optimization problem in Eq. (20), and $\beta\gg 1$. This imposes the exact parametrization of $A_{4}$, and we then successively decrease $\beta$ until one observes the emergence of spatial variations, and the corresponding error bars, see Fig. 2b (violet data). This ensures that our learning process is dominantly limited by shot noise, and not by having too few parameters in the ansatz.
@@ III.1.3 Learning the spatial variations of the couplings
III.2 Liouvillian Learning of a long-range interacting many-body system
In this section, we illustrate Liouvillian learning for a model system involving long-range spin-spin interactions, as realized in trapped-ion setups. The effective Hamiltonian is given by [9]
@@ III.2 Liouvillian Learning of a long-range interacting many-body system
$\mathcal{L}_{D}^{-}(\varrho)=\frac{\gamma^{-}}{2}\sum_{k=1}^{N}\left([\sigma_{k}^{-}\varrho,\sigma_{k}^{-,\dagger}]+[\sigma_{k}^{-},\varrho\sigma_{k}^{-,\dagger}]\right).$ (30)
Moreover, we assume the presence of dephasing terms originating from the presence of a fluctuating, classical magnetic field, that leads to shifts of the energy levels of the $k$-th ion proportional to $B_{k}^{\mathrm{cl}}(t)$. It can be shown that in the white-noise limit, i.e., with correlation function $\langle B_{k}^{\mathrm{cl}}(t)B_{l}^{\mathrm{cl}}(t^{\prime})\rangle=\Gamma_{kl}^{z}\delta(t-t^{\prime})$, this leads to Lindblad terms of the form
$\begin{split}\mathcal{L}_{D}^{z}(\varrho)=\sum_{k,l=1}^{N}\Gamma_{kl}^{z}\Big{\{}\sigma_{k}^{z}\varrho\sigma_{l}^{z}-\frac{1}{2}(\sigma_{k}^{z}\sigma_{j}^{z}\varrho+\varrho\sigma_{k}^{z}\sigma_{j}^{z})\Big{\}}.\end{split}$ (31)
@@
Figure 3: Hamiltonian learning in presence of collective dissipation. (a) The ratio $\tilde{\lambda}_{1}/\tilde{\lambda}_{2}$ after excluding the conserved quantity $Q$ for different regularization parameters $\beta$. The colored triangles describe the asymptotic values at $N_{\mathrm{runs}}\to\infty$. (b) Snapshots of the learned Hamiltonian parameters as indicated by the asterisks in the left panel. All error bars are computed via bootstrapping as explained in Appendix C, using $10$ samples. The integrals in Eq. (8) are evaluated using $B_{z}\Delta_{T}=1/128$, i.e., using $128$ time-steps, where at each time we use only a fraction of $1/128$ number of measurements per basis, compared to the quench times $B_{z}T\in[0,0.5,1]$. The auxiliary parameter $\xi$ for the additional constraints as introduced in Eq. (11) is set to $\xi=1000$. Model parameters: For the Hamiltonian we choose parameters as specified in Eq. (29) with $B_{z}=100$, $J_{0}=6/5B_{z}$ and $\alpha=1.5$. Moreover, in the Liouvillian, we choose $\gamma_{-}=B_{z}/20$ in Eq. (30), as well as $\Gamma_{kl}^{z}=\gamma_{k}^{z}\delta_{kl}+\Gamma_{0}$ with $\Gamma_{0}=B_{z}/40$ and $\gamma_{k}^{z}=3\Gamma_{0}$ in Eq. (31).
III.2.1 Ansatz for learning $H_{\mathrm{XY}}$
@@
In Section III.1, we have already illustrated our strategy for learning the operator content, as well as the relevant parameters of the Hamiltonian from minimal prior knowledge. Therefore, we start here with a parametrized ansatz, that already incorporates the algebraically decaying spin-spin interactions in $H_{\mathrm{XY}}$, that one would expect to find in a trapped-ion experiment, as well as a constant magnetic field. We choose as an ansatz for the Hamiltonian
$A(c_{0},\alpha,b)=\sum_{i<j}\frac{c_{0}}{|i-j|^{\alpha}}(\sigma_{i}^{x}\sigma_{j}^{x}+\sigma_{i}^{y}\sigma_{j}^{y})+b\sum_{k}\sigma_{k}^{z},$ (33)
@@
which depends non-linearly on the parameter $\alpha$, and hence, will require non-linear minimization of $\lambda_{1}$ over $\alpha$.
As we also want to learn the dissipative processes, we will start with an ansatz for the Liouvillian, comprising single-qubit Lindblad operators that typically appear in the context of trapped-ion experiments, i.e.,
$\mathcal{D}^{\mathrm{loc}}(\varrho)=\sum_{\eta\in\{+,-,z\}}\frac{d_{\eta}}{2}\sum_{k=1}^{N}\quantity([\sigma_{k}^{\eta}\varrho,\sigma_{k}^{\eta,\dagger}]+\mathrm{h.c.}),$ (34)
III.2.2 Learning a disordered Hamiltonian
Let us now illustrate learning of the long-range Hamiltonian $H_{\mathrm{XY}}$ using the physically motivated ansätze in Eqs. (33) and (34). Firstly, one notices that the ansatz $A$ contains the total magnetization defined by
$Q=B\sum_{k}\sigma^{z}_{k},$ (35)
In Fig. 3a we monitor $\tilde{\lambda}_{1}/\tilde{\lambda}_{2}$ as a function of $N_{\mathrm{runs}}$ for the ansatz $A_{XY}$, and for different values of the regularization parameter $\beta$. One starts with a large value of the regularization parameter, here, $\beta=100$, which strongly imposes the parametrization of $A_{XY}$, leading to reconstructed parameters without spatial disorder, see Fig. 3b for $N_{\mathrm{runs}}=2.1\times 10^{5}$ (blue data). Then further increasing $N_{\mathrm{runs}}$ reduces the size of the error bars in $\bm{c}^{\mathrm{rec}}$, as is also shown in Fig. 3b for $N_{\mathrm{runs}}=1.7\times 10^{6}$ (blue data). However, an ansatz with $\beta=100$ cannot account for disorder in the coupling terms in $H_{XY}$, which is shown by the plateau in $\tilde{\lambda}_{1}/\tilde{\lambda}_{2}$, here, starting at around $N_{\mathrm{runs}}\approx 10^{7}$ (blue line), which, again, is above our measurement budget. Nevertheless, we decrease the regularization parameter $\beta$, which leads to larger error bars, but also allows to learn some of the spatial disorder in $H_{\mathrm{XY}}$, as is shown in Fig. 3b for $\beta=2$ and $N_{\mathrm{runs}}=1.7\times 10^{6}$ (orange data). Note that already before reaching a plateau for $\beta=2$ in Fig. 3a (orange line) the error bars of $\bm{c}^{\mathrm{rec}}$ become very small, which can be seen in Fig. 3b for $\beta=2$ and $N_{\mathrm{runs}}=6.6\times 10^{6}$ (orange data). This suggest to further reduce $\beta$. Ultimately, this process of subsequently reducing $\beta$ and increasing $N_{\mathrm{runs}}$ allows us to learn $H_{\mathrm{XY}}$ up to statistical errors, and leads to $\tilde{\lambda}_{1}/\tilde{\lambda}_{2}\sim N_{\mathrm{runs}}^{-1/2}$ for the entire range of $N_{\mathrm{runs}}$.
$[M_{D}(\bm{d})]_{\kappa,(i,j)}=\sum_{k,l}d_{kl}[M^{(k,l)}]_{\xi,(i,j)}\\ =\int_{0}^{T}\sum_{k,l}d_{kl}\expectationvalue*{\sigma_{k}^{z}[h_{ij},\sigma_{l}^{z}]+\mathrm{h.c.}}_{\kappa,t}\mathrm{d}t,$ (36)
III.2.3 Learning collective dissipation
Figure 4: Learning collective dissipation using the additional constraints defined in the main text, and regularization strength $\beta=2$. (a) The error of the additional constraints $\Delta^{\mathrm{add}}$ as introduced in Eq. (38) for different ansätze for the dissipation as described the main text. (b) Snapshots of the learned dissipation rates $d_{k}^{\mathrm{rec}}$ as marked by the asterisks in the left panel. All error bars are computed via Bootstrapping as explained in Appendix C, using $10$ samples. Integrals for dissipation correction are evaluated as for Fig. 3.We further restrict $d_{-},d_{z},d_{z}^{\mathrm{col}}\geq 0$ and numerically search for the optimal rates within the bounds.

### v2

@@ III.2 Learning from generalized energy conservation
To begin, we illustrate Hamiltonian and Liouvillian learning from generalized energy conservation based on Eq. (11). We assume that the analog quantum simulator is governed by a master equation of the following form
$\frac{\mathrm{d}}{\mathrm{d}t}\varrho=-i[H,\varrho]+\sum_{\eta\in\{+,-,z\}}\frac{\gamma_{\eta}}{2}\sum_{k=1}^{N}\quantity([\sigma_{k}^{\eta}\varrho,\sigma_{k}^{\eta,\dagger}]+\mathrm{h.c.}).$ (16)
@@
1. Identifying the dominant terms of the Hamiltonian.
@@
leads to a much smaller value of $\lambda_{1}/\lambda_{2}$. We emphasize that learning with the ansatz $A_{2}$ only requires measurements of $\sigma_{x}^{\otimes N}$ and $\sigma_{z}^{\otimes N}$. Therefore, some of the measurements performed for $A_{1}$ can be recycled for learning the parameters of $A_{2}$. By further measuring, a plateau in $\lambda_{1}/\lambda_{2}$ starts to emerge at $N_{\mathrm{runs}}\approx 3\times 10^{5}$, as shown in Fig. 3a (orange line). At this point we have identified all the dominant terms in $H$. To continue the learning process we need to extend the operator ansatz. Therefore, in the next step we want to learn subdominant terms of $H$, as well as the Liouvillian.
2. Learning the sub-dominant terms of the Hamiltonian, and learning the Liouvillian.
@@
3. Learning the spatial variations of the couplings.
As can be seen in Fig. 3a (red line), the plateau in $A_{4}$ is reached at $N_{\mathrm{runs}}\approx 10^{7}$, which again exceeds our available measurement budget. Therefore, we are not allowed to conclude that ansatz $A_{4}$ is insufficient to describe the Hamiltonian in Eq. (II.1) by only considering the ratio $\lambda_{1}/\lambda_{2}$. However, at this point, one can choose to test other ansätze and compare their corresponding reconstructed parameters, or to see if smaller values of $\lambda_{1}/\lambda_{2}$ are achieved. As one typically expects small spatial variations in the couplings of $H$, and $A_{4}$ was already a good approximation of $H$, we want to test for spatial variations on top of the spatially homogeneous ansatz $A_{4}$. To this end, we simply treat the parametrization of the ansatz $A_{4}$ as a regularization of the cost function, see Appendix C for technical details. That is, we start with the optimization problem in Eq. (14), and add a penalty term with control parameter $\beta$ for deviating from a given parametrization. Choosing $\beta\gg 1$ imposes the exact parametrization of $A_{4}$, and we then successively decrease $\beta$ until one observes the emergence of spatial variations, and the corresponding error bars, see Fig. 3b (purple data). This ensures that our learning process is dominantly limited by shot noise, and not by having too few parameters in the ansatz.
@@
Figure 4: Learning collective dissipation using random measurements. (a) The ratio $\lambda_{1}/\lambda_{2}$ of the matrix $(M,-b)$ for different regularization parameters $\beta$. (b,c) Snapshots of the learned Hamiltonian parameters, $c^{\rm rec}_{j}$, corresponding to the parameters $J_{kl}$ and $B_{i}$ defined in Eq. (23). The required measurement budget is also indicated in the left panel. Black crosses correspond to the parameters of the model Hamiltonian specified below. All error bars are computed via bootstrapping as explained in Appendix D, using $40$ samples. The time-traces in Eq. (13) are evaluated using $B_{z}\Delta_{T}=1/20$, i.e., using $20$ time-steps, using the same amount of random measurements at each time up to a time of $B_{z}T=0.5$. Expectation values at $t=0$ are calculated exactly from the initial state. Model parameters: For the Hamiltonian we choose parameters as specified in Eq. (24) with $B_{z}=100$, $J_{0}=6/5B_{z}$ and $\alpha=1.5$. Moreover, in the Liouvillian, we choose $\gamma_{-}=B_{z}/20$ in Eq. (25), as well as $\Gamma_{kl}^{z}=\gamma_{k}^{z}\delta_{kl}+\Gamma_{0}$ with $\Gamma_{0}=B_{z}/40$ and $\gamma_{k}^{z}=3\Gamma_{0}$ in Eq. (26).
III.3 Learning from generalized Ehrenfest theorem
In this section, we will learn the Hamiltonian and Liouvillian of a model system from the generalized Ehrenfest theorem in Eq. (7). As this method requires the estimation of expectation values of many observables, $\{O_{i}\}$, we will use the randomized measurement toolbox [47] to obtain the necessary data. This will also allow us to demonstrate Hamiltonian learning in large systems of tens of qubits, and provide evidence that given a fixed number of shots the reconstruction error is independent of the system size, if the number of Hamiltonian parameters can be kept constant.
We consider a model system involving long-range spin-spin interactions, as realized in trapped-ion setups. The effective Hamiltonian is given by [9]
@@ III.3 Learning from generalized Ehrenfest theorem
$\mathcal{L}^{-}(\varrho)=\frac{\gamma^{-}}{2}\sum_{k=1}^{N}\left([\sigma_{k}^{-}\varrho,\sigma_{k}^{-,\dagger}]+[\sigma_{k}^{-},\varrho\sigma_{k}^{-,\dagger}]\right).$ (25)
Moreover, we consider the presence of dephasing terms originating from the presence of a fluctuating, classical magnetic field, that leads to shifts of the energy levels of the $k$-th ion proportional to $B_{k}^{\mathrm{cl}}(t)$. It can be shown that in the white-noise limit, i.e., with correlation function $\langle B_{k}^{\mathrm{cl}}(t)B_{l}^{\mathrm{cl}}(t^{\prime})\rangle=\Gamma_{kl}^{z}\delta(t-t^{\prime})$, this leads to Lindblad terms of the form
$\begin{split}\mathcal{L}^{z}(\varrho)=\sum_{k,l=1}^{N}\Gamma_{kl}^{z}\Big{\{}\sigma_{k}^{z}\varrho\sigma_{l}^{z}-\frac{1}{2}(\sigma_{k}^{z}\sigma_{j}^{z}\varrho+\varrho\sigma_{k}^{z}\sigma_{j}^{z})\Big{\}}.\end{split}$ (26)
@@
1. Hamiltonian and Liouvillian learning using randomized measurements.
@@
In Section III.2, we already illustrated learning of the operator content, and the relevant parameters of the Hamiltonian from minimal assumptions. Here, we want to focus on distinguishing local from collective dephasing, and analyse the scaling of the reconstruction error. Therefore, we start with a parametrized ansatz, that already incorporates the algebraically decaying spin-spin interactions in $H_{\mathrm{XY}}$, that one would expect to find in a trapped-ion experiment, as well as a constant magnetic field. We choose as an ansatz for the Hamiltonian
$A(\bm{c})=\sum_{i<j}\frac{c_{0}}{|i-j|^{\alpha}}(\sigma_{i}^{x}\sigma_{j}^{x}+\sigma_{i}^{y}\sigma_{j}^{y})+b\sum_{k}\sigma_{k}^{z},$ (28)
@@
which depends non-linearly on the parameter $\alpha$, and hence, will require non-linear minimization of $\lambda_{1}$ over $\alpha$ as explained in Appendix C. As an ansatz for the Liouvillian we choose
$\mathcal{D}(\varrho)=\frac{d_{-}}{2}\sum_{k=1}^{N}\quantity([\sigma_{k}^{-}\varrho,\sigma_{k}^{-,\dagger}]+\mathrm{h.c.})+\sum_{k,l=1}^{N}d^{z}_{|k-l|}\mathcal{D}^{z}_{kl}(\varrho),$ (29)
that includes single-spin spontaneous decay with jump operator $\sigma_{k}^{-}$, and single-spin as well as multi-spin collective dephasing $\mathcal{D}^{z}_{kl}$ with operator structure as described in Eq. (26).
To obtain constraint operators, $\{O_{i}\}$, we proceed as follows; we choose a set of random Pauli measurements, $\{P_{\bm{\alpha}}\}=\{\sigma^{\alpha_{1}}\otimes\dots\otimes\sigma^{\alpha_{N}}\}$, with $\alpha_{i}\in\{x,y,z\}$, which we measure at consecutive times from $t=0$ up to $t=T$, similar as before (see Fig. 4 for details). From this set of Pauli measurements we extract all one- and two-qubit operators that are contained in $\{P_{\bm{\alpha}}\}$, as well as their commutators with the ansatz operators and jump operators (in case those are contained). For these operators and commutators we extract expectation values to obtain the data necessary to estimate the constraint matrix in Eq. (7). Choosing the set $\{P_{\bm{\alpha}}\}$ large enough ensures that the necessary operators are contained in this set.
The results are presented in Fig. 4. In Fig. 4(a) we show the ratio of the two smallest singular values of the matrix $(K_{H},K_{D},-\bm{b})$, which can be obtained by rewriting Eq. (7), where $K_{H},K_{D}$, and $\bm{b}$ are the constraint matrices and vector respectively. The quantity $\lambda_{1}/\lambda_{2}$ shows a similar scaling behavior in early and late stages of the learning procedure as the analog quantity in Section III.2. Moreover, it correctly recognizes the insufficiency of the translation invariant ansatz ($\beta=0.1$, blue), as well as the larger number of parameters of the almost unparametrized ansatz ($\beta=0.001$, orange). In Fig. 4(b) we show the learned Hamiltonian parameters (top) and dissipation rates (bottom) for both ansätze, and a total measurement budget of $N_{\mathrm{runs}}=2.6\times 10^{6}$, in comparison to the true values (black). For the translation invariant ansatz (blue) one notices that while the reconstructed parameters of the Hamiltonian roughly resemble the true parameters, the dissipation rates turn out wrong. We have numerically verified that this artefact does not disappear for larger measurement budget. However, by lowering the penalty $\beta$ for deviating from translation inference the solution $\bm{c}^{\rm rec}$ correctly accounts for disorder, and the dissipation rates are learned accurately.
2. Comparison between learning form Ehrenfest constraints and energy conservation.
To make a fair comparison between Hamiltonian and Liouvillian learning from energy conservation and the Ehrenfest theorem, we have learned the same model system using Eq. (11). This, however, also requires the measurement of additional constraints to learn the overall scale, and to distinguish local from collective dephasing due to symmetries of $H_{XY}$. The results are discussed in Appendix F. One observes, that in a system with conserved quantity, learning from the Ehrenfest theorem in Eq. (7) requires less measurements, when compared to learning from generalized energy conservation in Eq. (11). For instance, verifying the presence of collective dephasing using the Ehrenfest theorem requires $N_{\rm runs}\approx 10^{6}$ (cf. Figure 4), compared to $N_{\rm runs}\approx 10^{7}$ when using energy conservation (cf. Figure 9 in Appendix F).
Figure 5: Hamiltonian learning with random measurements using a fixed total number of random measurements $N_{\mathrm{runs}}=1.9\times 10^{6}$. The integrals in Eq. (13) are solved using $\Delta_{T}=T/10$, i.e., using $10$ equidistant time-steps, where at each time we use the same amount of measurements of $N_{\rm bases}=2000$ random bases, with a total time of $B_{z}T=0.1$ for $B_{z}$ as in Fig. 8. Expectation values at $t=0$ are calculated exactly from the initial state. (a) The error in the learned parameters as function of system size for different regularization strengths $\beta$. (b) Learned parameters for different system sizes $N\in[10,20,40,60]$ for a fixed regularization strength $\beta=0.001$. Model parameters: We choose the same interaction strengths $J_{ij}$ as in Fig. 8, but setting the dissipation to zero and $B_{z}=0$.
3. Scaling to larger system sizes.
We now want to study the scaling behavior of the reconstruction error of the Hamiltonian parameters, by varying system size from $N=10$ up to $N=60$ spins, for a Hamiltonian that is dominantly described by a constant number of parameters. In order to simplify the presentation and to reduce the classical computation time needed for simulating the measurements we make the following modifications to our model system: (i) We set the magnetic field and the dissipation rates to zero. (ii) We consider a Hamiltonian of the form
$H=\sum_{k=1}^{N_{\mathrm{sub}}}\sum_{i\neq j=1}^{N}J_{ij}\left(\sigma_{(kN+i)}^{x}\sigma_{(kN+j)}^{x}+\sigma_{(kN+i)}^{y}\sigma_{(kN+j)}^{y}\right),$ (30)
consisting of $N_{\rm sub}$ non-interacting subsystems of $N=10$ spins each, which makes it easy to simulate on a classical computer.
To learn the Hamiltonian in Eq. (30) for the different system sizes, we want to compare leaning without parametrization, where the number of parameters in the ansatz grows with systems size, to learning using a parametrized ansatz where the number of parameters is kept constant with system size. Therefore, we choose an ansatz
$A(\bm{c})=\sum_{k=1}^{N_{\mathrm{sub}}}\sum_{i\neq j=1}^{10}\frac{c_{0}}{|i-j|^{\alpha}}\left(\sigma_{i+k}^{x}\sigma_{j+k}^{x}+\sigma_{i+k}^{y}\sigma_{j+k}^{y}\right)\\ +\sum_{k\neq l=1}^{N_{\mathrm{sub}}}\sum_{i\neq j=1}^{10}c_{1}\left(\sigma_{i+k}^{x}\sigma_{j+l}^{x}+\sigma_{i+k}^{y}\sigma_{j+l}^{y}\right),$ (31)
with a constant number of parameters, and that depends non-linearly on the parameter $\alpha$ similar to Eq. (28). As in the previous example the parametrization of this ansatz is not strictly enforced, but depends on the chosen regularization strength $\beta$ (see Appendix C for technical details). We emphasize that for $\beta=0$ we are learning from an unparametrized ansatz, with an independent parameter for each operator in the ansatz. In this case, our reconstructed Hamiltonian does not have the subsystem-structure of the true Hamiltonian in Eq. (30) inherently built-in; rather, this structure is learned by including all possible inter-subsystem interaction terms.
We learn the Hamiltonian in Eq. (30) for $N=10,\dots,60$ systems. We choose a total of $2000$ random Pauli measurements for each of a total of $11$ time-steps up to $T=1$, leading to a total measurement budget of $N_{\rm runs}=1.5\times 10^{6}$. For the constraint matrix $M$ and vector $b$ in Eq. (10) we choose the constraints to be all possible one- and two-qubit constraint operators, and evaluate the corresponding constraint matrix $M$ and vector $b$. We then determine the learned coefficients $\bm{c}^{\mathrm{rec}}$ from Eq. (10).
The result is shown in Fig. 5. In Fig. 5(a) we show the relative error of the learned parameters defined by
$\Delta\bm{c}=\|\bm{c}^{\mathrm{rec}}-\bm{c}^{\mathrm{sim}}\|/\|\bm{c}^{\mathrm{sim}}\|$ (32)
as a function of the system size $N$. We find that for weakly, or unparamatrized ansätze, for which the number of learned parameters grows with system size, the reconstruction error grows only moderately. This is likely explained by the fact that in larger systems a single round of measurements extracts more information as more correlation functions can be measured in a single shot. Moreover, for parametrized ansätze with a system-size independent number of parameters the reconstruction error also remains independent of the system-size.